Mai, Ali, and Ryan sent a total of 139 text messages over their cell phones during the weekend. Ryan sent 4 times as many messages as Mai. Mai sent 5 more messages than Ali. How many messages did they each send?
Answer by kingme18(98)

Let's call the number of texts Mai sent m. Ryan sent 4*m (4 times as many as Mai). Mai sent 5 more than Ali, so Ali sent 5 less than Mai: m - 5
Add the number they each sent to get the total: 139
(m) + (4*m) + (m - 5) = 139
6*m - 5 = 139
6*m=144
m = 24
Mai sent 24. This means Ryan sent 4*24=96 messages, and Ali sent 24-5=19 messages.
Check: 24 + 96 + 19 = 139